Use the function to evaluate the indicated expressions and simplify.
f(x) = 2x^2 + 3
f(x+2) = (blank)
f(x) + f(2) = (blank)

Found 2 solutions by josgarithmetic, math_tutor2020:
Answer by josgarithmetic(39617)

Those appear stated in a non-standard way.

Answer by math_tutor2020(3816)


f(x) = 2x^2 + 3
f(x+2) = 2(x+2)^2 + 3 ..... every x replaced with x+2
f(x+2) = 2(x^2+4x+4) + 3 .... FOIL rule
f(x+2) = 2x^2+8x+8 + 3
f(x+2) = 2x^2+8x+11

f(x) = 2x^2+3
f(2) = 2*2^2+3 .... every x replaced with 2
f(2) = 2*4+3
f(2) = 8+3
f(2) = 11
f(x)+f(2) = (2x^2+3) + (11)
f(x)+f(2) = 2x^2+14


Extra info:
We can see that
f(x+2) = f(x)+f(2)
is not a true equation since we ended up with different polynomial results.
In more general terms, f(a+b) = f(a)+f(b) is also not true.